# DeTracker: Motion-decoupled Vehicle Detection and Tracking in Unstabilized Satellite Videos

### v1

@@ IV-A Framework Overview
Figure 3: Overview of the proposed DeTracker. The framework consists of three components: (1) a feature extraction backbone for acquiring multi-scale spatial representations from input frames; (2) a Global–Local Motion Decoupling (GLMD) module for disentangling platform-induced global motion from target-level local motion; (3) a Temporal Dependency Feature Pyramid (TDFP) module for aggregating temporal information and modeling cross-frame dependencies.
@@
Unlike conventional MOT pipelines that rely on stabilized input or post-hoc association, the proposed DeTracker directly operates on unstabilized satellite video and performs detection and tracking in an end-to-end manner. This design avoids hand-crafted motion compensation and enables the model to learn intrinsic motion structures directly from data, rather than relying on optical flow or Kalman-based heuristics.
As illustrated in Fig. 3, DeTracker consists of three components: a feature extraction backbone, a GLMD module, and a TDFP module. Specifically, given a video sequence $V=\{{I^{n}\mid n=1,2,\ldots,N}\}$, where $I^{n}$ is the $n$-th frame and $N$ is the total number of frames, we employ DLA34 [46] as the backbone and take a contiguous clip of $k$ frames ending at time $t$, i.e., $V_{clip}=\{I^{\tau}\mid\tau=t-k+1,\ldots,t\}$, as input. For each input frame $I_{\tau}$, the backbone outputs a set of multi-scale feature maps $\{f_{i}^{\tau}\mid i=1,2,3,4\}$, where $i$ indexes the feature level from low to high dimension.
Subsequently, mid-level features are selected for motion modeling, as shallow features susceptible to noise, while deep features tend to collapse tiny objects into single pixels. The use of a stable $k$-frame input allows motion cues to accumulate in feature space instead of being reconstructed through explicit data association. Specifically, the mid-level features $\{f^{\tau}_{j}\mid j=2,3\}$ from adjacent frames are fed into the GLMD module to correct misalignment in the current frame, where global translation trends and local fine-grained offsets are modeled, respectively. Together with low-level and high-level features, the motion-refined mid-level features form a updated feature set $\{\hat{f}_{i}^{\tau}=(f_{1}^{\tau},\hat{f}_{2}^{\tau},\hat{f}_{3}^{\tau},f_{4}^{\tau})\mid\tau=t-k+1,\dots,t\}$. Subsequently, $\{\hat{f}_{i}^{\tau}\}$ is fed into the TDPF module, yielding a compact, low-dimensional feature set $\{F^{\tau}\mid\tau=t-k+1,\ldots,t\}$ enriched with semantic and temporal information. Finally, $\{F^{\tau}\}$ is fed into the detection head, where object detection and tracking are jointly performed via center-point prediction.
IV-B Global-Local Motion Decoupling
As illustrated in Fig. 4, the proposed GLMD module adopts a dual-branch architecture: a global alignment branch that learns affine transformations to achieve coarse global registration, and a local refinement branch that further models local motion via adaptive, deformable sampling.
@@ IV-B1 Global Alignment
$Atten_{FCA}=softmax(\frac{Q_{t}K_{t-1}^{T}}{\sqrt{d}})V_{t-1},$ (5)
$Atten_{BCA}=softmax(\frac{Q_{t-1}K_{t}^{T}}{\sqrt{d}})V_{t},$ (6)
@@
$Atten=softmax(\frac{Q_{h}K_{h}^{T}}{\sqrt{d}})V_{h},$ (8)
@@ IV-B2 Local Refinement
Next, the global transformation matrix $R_{t}$ is decomposed into a linear component $A\in\mathbb{R}^{2\times 2}$ and a translation vector $b\in\mathbb{R}^{2}$. To incorporate the affine alignment results into local alignment, we construct a fundamental offset term $\Delta{P_{k}}$ for each convolution kernel position $P_{k}$ in feature space, To accommodate variations in feature resolution, we perform spatial scale normalization using a diagonal matrix: $D=diag(W/2,H/2)$, where $W$ and $H$ represent the image width and height, respectively. The resulting affine prior offset $\Delta{P_{k}}$ provide a principled initialization for deformable sampling, upon which the local branch learns fine-grained corrections:
@@
Finally, the motion-aware feature map $r^{t}_{j}$, the totle offset $\mathcal{O}$, and the modulation mask $m$ are fed into a deformable convolution layer. The resulting features are normalized by batch normalization to obtain the refined feature map $\{\hat{f}_{j}^{t}\mid j=2,3\}$, which incorporates both global alignment and local refinement. This process is expressed as:
$\mathcal{O}=\Delta{P_{k}}+\Delta{P_{k}^{res}}$ (13)
$F^{\tau}_{j}=SL(BN(De(r^{\tau}_{j},\mathcal{O},m)))$ (14)
@@ IV-C Temporal Dependency Feature Pyramid
Temporal dependency modeling is crucial for tiny-object tracking in satellite videos, as motion cues can be accumulated across multiple consecutive frames to overcome severe noise, weak appearance, and frequent occlusions. Among existing temporal modeling strategies, ConvGRU [1] provides a more appropriate trade-off between robustness and efficiency compared to optical flow–based methods and Transformer-based architectures. Optical flow estimation becomes unreliable when inter-frame displacements fall below two pixels, which is common for distant vehicles in satellite imagery. Transformer-based models, although effective for long-range dependency modeling, incur substantial computational overhead when applied to video sequences. In contrast, ConvGRU incrementally aggregates local motion evidence in a recurrent manner, preserving identity continuity while maintaining low computational cost. Therefore, we here adopt ConvGRU to efficiently model temporal dependencies.
As illustrated in the lower part of Fig. 3, the feature sequence $\{\hat{f}_{i}^{\tau}=(f_{1}^{\tau},\hat{f}_{2}^{\tau},\hat{f}_{3}^{\tau},f_{4}^{\tau})\mid\tau=t-k+1,\dots,t\}$ is first processed by ConvGRU to obtain a temporally fused feature set $\{\tilde{f}_{i}^{\tau}\}$. In this process, for the input $\{\hat{f}_{i}^{t}\mid i=1,2,3,4\}$ in the sequence and the hidden state $\{\hat{f}_{i}^{t-1}\mid i=1,2,3,4\}$ from the previous frame, ConvGRU computes the update and reset gates followed by the candidate state and the new hidden state:
$z$ $=\sigma{[W_{z}(F_{i}^{\tau},h)]},$ (15)
$r$ $=\sigma{[W_{r}(F_{i}^{\tau},h]},$
$\tilde{h}$ $=tanh\{\sigma{[W_{h}(F_{i}^{\tau},r\odot{h})]}\},$
$h$ $=(1-z)\odot{h}+z\odot{\tilde{h}}$
@@
Finally, to capture the dynamic evolution of objects in a higher-level semantic space, we apply a second ConvGRU aggregation over the fused feature sequence $\{F^{\tau}\mid\tau=t-k+1,\ldots,t\}$. This temporal aggregation further stabilizes object representations and strengthens identity continuity across frames. The resulting features are then used to generate the heatmap, width–height regression, and offset regression branches, enabling joint object detection and tracking.
@@ IV-D Objective Functions
$\mathcal{L}_{warp}=\mathbb{E}_{p\in\Omega}[1-\frac{\langle f_{warp}(p),f_{re}(p)\rangle}{\parallel{f_{warp}(p)}\parallel{}_{2}\times\parallel{f_{re}(p)\parallel{}_{2}}+\epsilon}],$ (16)
where $\langle\cdot\rangle$ denotes the inner product, $\Omega$ represents the spatial domain of the feature map, and $\epsilon$ is a small constant for numerical stability. This cosine similarity–based formulation encourages aligned features to remain semantically consistent while being invariant to feature magnitude.
@@ V-A1 Datasets
The proposed method is evaluated on 1) simulated unstabilized dataset SDM-Car-SU, which covers diverse background motion patterns with varying speeds (1–3 pixels per frame) and motion directions, enabling systematic assessment under different levels of unstabilized motion; and 2) real unstabilized satellite videos acquired by the Luojia-3 01 satellite to validate its effectiveness in practical scenarios.
@@ V-A2 Baselines
We compare the proposed DeTracker with a range of state-of-the-art trackers, including TBD-based approachs, including DeepSORT [40], CKDNet-SMTNet [12] and ByetTrack [52], and JDT-based approachs, including CenterTrack [60], FairMOT [53], DSFNet [42], MP2Net [56], DAF [47], PIFTrack [5] and MOSAIC-Tracker [63]. To ensure fair comparisons, all methods are trained and evaluated on the same server, with hyperparameters and training settings are adopted from the original papers or official open-source implementations.
@@ V-A3 Evaluation Metrics
For satellite video multi-object tracking, we adopt MOTA to measure overall tracking performance, with a primary emphasis on detection accuracy, and IDF1 to evaluate identity retention capability, which reflects the quality of track association:
$\mathrm{MOTA}=1-\frac{FN+FP-ID_{S}}{GT},$ (18)
@@
where $FP$, FNFN, and $ID_{S}$ denote false positives, false negatives and identity switches, respectively, $GT$ is the number of ground-truth objects, $ID_{TP}$, $ID_{FP}$, and $ID_{FN}$ are the numbers of true-positive, false-positive, and false-negative identity matches. In addition, we report identity precision and recall (IDP and IDR), as well as standard MOT statistics including Precision (Prcn), Recall (Rcll), Mostly Tracked (MT), and Mostly Lost (ML), which provide complementary insight into tracking quality and robustness.
@@ V-A4 Implementation Details
Prior to training, standard data augmentation techniques were applied, including inversion, random flipping, and rotation. The initial learning rate was set to $1.25\times 10^{-4}$ and decayed using a cosine annealing schedule. The model was trained for 150 epochs with a batch size of 4. The loss weghting parameters $\alpha$ and $\beta$ were set to 0.2 and 0.3, respectively. The input frame number $k$ was set to 5. All experiments were conducted on a single NVIDIA A100 Tensor Core GPU.
V-B Comparison with the State-of-the-Arts
@@ V-B1 Quantitative Analysis
As shown in Table I, we quantitatively compare DeTracker with the baseline methods on the SDM-Car-SU dataset. It can be observed that DeTracker consistently achieves the best overall performance across all motion intensity subsets (U1–U3), where U1–U3 correspond to increasing platform motion intensities from 1 to 3 pixels per frame. Specifically, in terms of the two core metrics, DeTracker achieves 61.1%, 55.3%, 52.4% in MOTA and 68.7%, 67.2%, 63.8% in IDF1 on U1, U2, and U3, respectively, outperforming all competing methods by a clear margin. Compared with the second-best method, PIFTrack [5], DeTracker yields an average improvement of 5.1% in MOTA and 2.9% in IDF1 over three motion intensities, clearly demonstrating the effectiveness of our MOT method in instabilized videos. Furthermore, DeTracker substantial reduces error-related metrics, such as IDS, FP, and FN, indicating stronger robustness in both data association and improved recall of tiny objects.
To further evaluate generalization under real on-orbit conditions, we conduct experiments on unstabilized real satellite videos. As reported in Table II, DeTracker achieves 45.3% in MOTA and 67.8% in IDF1, surpassing PIFTrack by 8.3% and 7.7%, respectively. These results confirm that DeTracker not only performs strongly on simulated data but also maintains maintains high accuracy and robustness in real unstabilized satellite scenarios.
@@
FarMOT [53] DSFNet [42] MP2Net [56] PIFTrack [5] DeTracker (Ours) Ground truth
Figure 5: Examples of detection results for vehicles moving in different directions. From the first line to the third line are frames 50, 120, and 189 respectively and vehicles undergo lateral and longitudinal motion. Green boxes denote true positives and correct predictions, red boxes indicate false positives and yellow boxes represent false negatives.
@@
Input frame FarMOT [53] DSFNet [42] MP2Net [56] PIFTrack [5] DeTracker (Ours) Ground truth
@@ V-B2 Qualitative Analysis
We visualize detection results on randomized camera-shake video sequences with diverse vehicle motion directions for different methods, as illustrated in Fig. 5. Under complex backgrounds and severe platform jitter, most baseline methods suffer from frequent false detections and missed objects, particularly in regions with strong illumination changes, shadows, and road boundaries, where bounding boxes often drift or become misaligned. In contrast, DeTracker accurately localizes vehicle targets across consecutive frames and reduces both false positives and missed detections (see the red and yellow highlighted regions).
Fig. 6 further visualizes trajectory predictions from different trackers, where all trajectories are projected onto the last frame of each sequence and color coded by vehicle identity. As shown in the first row, DeTracker generates smooth and continuous trajectories that closely follow the ground truth paths, remain well aligned with road geometry, and exhibit minimal jitter. In contrast, most baselines suffer from notable trajectory fluctuations at turning regions, with some trajectories drifting along background structures. In the second row, baseline methods produce fragmented or discontinuous trajectories, indicating identity switches and unstable temporal association. Notably, MP2Net and DSFNet exhibit clear trajectory offsets and breaks, whereas DeTracker consistently predicts complete and coherent trajectories. We attribute this performance gain to the GLMD module, which decouples global background motion and enables feature-level alignment, and the TDPF module, which strengthens cross-frame dependency modeling, thereby preserving trajectory coherence and identity consistency over long-term tracking.

### v2

@@ IV-A Framework Overview
Figure 4: Overview of the proposed DeTracker. The framework consists of three components: (1) a feature extraction backbone for acquiring multi-scale spatial representations from input frames; (2) a Global–Local Motion Decoupling (GLMD) module for disentangling platform-induced global motion from target-level local motion; and (3) a Temporal Dependency Feature Pyramid (TDFP) module for aggregating temporal information and modeling cross-frame dependencies.
Unlike conventional MOT pipelines that rely on stabilized inputs or post-hoc association, the proposed DeTracker directly operates on unstabilized satellite video and performs detection and tracking in an end-to-end manner. This design avoids hand-crafted motion compensation and enables the model to learn intrinsic motion structures directly from data, rather than relying on optical flow or Kalman-based heuristics.
As illustrated in Fig. 4, DeTracker consists of three components: a feature extraction backbone, a GLMD module, and a TDFP module. Specifically, given a video sequence $V=\{{I^{n}\mid n=1,2,\ldots,N}\}$, where $I^{n}$ is the $n$-th frame and $N$ is the total number of frames, we employ DLA34 [53] as the backbone and take a contiguous clip of $k$ frames ending at time $t$, i.e., $V_{clip}=\{I^{\tau}\mid\tau=t-k+1,\ldots,t\}$, as input. For each input frame $I^{\tau}$, the backbone outputs a set of multi-scale feature maps $\{f_{i}^{\tau}\mid i=1,2,3,4\}$, where $i$ indexes the feature level from low to high dimension.
Subsequently, mid-level features are selected for motion modeling, as shallow features susceptible to noise, while deep features tend to collapse tiny objects into single pixels. The use of a stable $k$-frame input allows motion cues to accumulate in feature space instead of being reconstructed through explicit data association. Specifically, the mid-level features $\{f^{\tau}_{j}\mid j=2,3\}$ from adjacent frames are fed into the GLMD module to correct misalignment in the current frame, where global translation trends and local fine-grained offsets are modeled, respectively. Together with low-level and high-level features, the motion-refined mid-level features form an updated feature set $\{\hat{f}_{i}^{\tau}=(f_{1}^{\tau},\hat{f}_{2}^{\tau},\hat{f}_{3}^{\tau},f_{4}^{\tau})\mid\tau=t-k+1,\dots,t\}$. Then, $\{\hat{f}_{i}^{\tau}\}$ is fed into the TDFP module, yielding a compact, low-dimensional feature set $\{F^{\tau}\mid\tau=t-k+1,\ldots,t\}$ enriched with semantic and temporal information. Finally, $\{F^{\tau}\}$ is fed into the detection head, where object detection and tracking are jointly performed via center-point prediction.
IV-B Global–Local Motion Decoupling
@@
As illustrated in Fig. 5, the proposed GLMD module adopts a dual-branch architecture: a global alignment branch that learns affine transformations in the feature space to suppress platform-dominated background motion, and a local refinement branch that models target-intrinsic motion via adaptive, deformable sampling.
@@ IV-B1 Global Alignment
$Atten_{FCA}$ $=\mathrm{softmax}\left(\frac{Q_{t}K_{t-1}^{T}}{\sqrt{d}}\right)V_{t-1},$ (5)
$Atten_{BCA}$ $=\mathrm{softmax}\left(\frac{Q_{t-1}K_{t}^{T}}{\sqrt{d}}\right)V_{t},$ (6)
@@
$Atten$ $=\mathrm{softmax}\left(\frac{Q_{h}K_{h}^{T}}{\sqrt{d}}\right)V_{h},$ (8)
@@ IV-B2 Local Refinement
Next, the global transformation matrix $R_{t}$ is decomposed into a linear component $A\in\mathbb{R}^{2\times 2}$ and a translation vector $b\in\mathbb{R}^{2}$. To incorporate the affine alignment results into local alignment, we construct a fundamental offset term $\Delta{P_{k}}$ for each convolution kernel position $P_{k}$ in feature space. To accommodate variations in feature resolution, we perform spatial scale normalization using a diagonal matrix: $D=diag(W/2,H/2)$, where $W$ and $H$ represent the image width and height, respectively. The resulting affine prior offset $\Delta{P_{k}}$ provides a principled initialization for deformable sampling, upon which the local branch learns fine-grained corrections:
@@
Finally, the motion-aware feature map $r^{t}_{j}$, the total offset $\mathcal{O}$, and the modulation mask $m$ are fed into a deformable convolution layer. The resulting features are normalized by batch normalization to obtain the refined feature map $\{\hat{f}_{j}^{t}\mid j=2,3\}$, which incorporates both global alignment and local refinement. This process is expressed as:
$\mathcal{O}$ $=\Delta{P_{k}}+\Delta{P_{k}^{res}},$ (13)
$\hat{f}_{j}^{t}$ $=SL(BN(De(r^{t}_{j},\mathcal{O},m))),$ (14)
@@ IV-C Temporal Dependency Feature Pyramid
Temporal dependency modeling is crucial for tiny-object tracking in satellite videos, as aggregating motion cues across consecutive frames mitigates noise, weak appearance features, and occlusions, especially under limited spatial resolution and subtle target motion. Among existing temporal modeling strategies, ConvGRU [1] provides a superior trade-off between robustness and computational efficiency compared to optical-flow-based and Transformer-based methods. When inter-frame displacement is smaller than two pixels and target motion is subtle, optical flow estimation often becomes unreliable. Moreover, optical-flow-based methods typically rely on explicit motion estimation via feature matching or cost volume construction, which introduces additional computational overhead. In contrast, although Transformer-based models are effective in capturing long-range dependencies, they usually incur high computational complexity when applied to video sequences. Therefore, we adopt ConvGRU to efficiently model temporal dependencies.
As illustrated in the lower part of Fig. 4, the feature sequence $\{\hat{f}_{i}^{\tau}=(f_{1}^{\tau},\hat{f}_{2}^{\tau},\hat{f}_{3}^{\tau},f_{4}^{\tau})\mid\tau=t-k+1,\dots,t\}$ is first processed by ConvGRU to obtain a temporally fused feature set $\{\tilde{f}_{i}^{\tau}\}$. Given the input $\{\hat{f}_{i}^{t}\mid i=1,2,3,4\}$ in the sequence and the hidden state $\{\hat{f}_{i}^{t-1}\mid i=1,2,3,4\}$ from the previous frame, ConvGRU computes the update and reset gates followed by the candidate state and the new hidden state:
$z$ $=\sigma{[W_{z}(\hat{f}_{i}^{\tau},h)]},$ (15)
$r$ $=\sigma{[W_{r}(\hat{f}_{i}^{\tau},h]},$
$\tilde{h}$ $=tanh\{\sigma{[W_{h}(\hat{f}_{i}^{\tau},r\odot{h})]}\},$
$h$ $=(1-z)\odot{h}+z\odot{\tilde{h}}$,
@@
Finally, to capture the dynamic evolution of objects in a higher-level semantic space, we apply a second ConvGRU aggregation over the fused feature sequence $\{F^{\tau}\mid\tau=t-k+1,\ldots,t\}$. This temporal aggregation establishes a cross-frame memory mechanism within the feature space, thereby implicitly reinforcing identity consistency modeling. It not only further stabilizes object representations but also significantly enhances the robustness of cross-frame associations. Subsequently, these features are utilized to generate the heatmap, width-height regression, and offset regression branches, enabling joint object detection and tracking.
@@ IV-D Objective Functions
$\mathcal{L}_{warp}=\mathbb{E}_{p\in\Omega}\left[1-\frac{\langle f_{warp}(p),f_{re}(p)\rangle}{\lVert f_{warp}(p)\rVert_{2}\lVert f_{re}(p)\rVert_{2}+\epsilon}\right],$ (16)
where $\langle\cdot\rangle$ denotes the inner product, $\Omega$ represents the spatial domain of the feature map, and $\epsilon$ is a small constant for numerical stability. This cosine similarity-based formulation encourages aligned features to remain semantically consistent while being invariant to feature magnitude.
@@ V-A1 Datasets
The proposed method is evaluated on two types of data: (i) simulated unstabilized dataset SDM-Car-SU, which covers diverse background motion patterns with varying speeds (1–3 pixels per frame) and motion directions, enabling systematic assessment under different levels of unstabilized motion; and (ii) real unstabilized satellite videos acquired by the Luojia-3 01 satellite to validate its effectiveness in practical scenarios.
@@ V-A2 Baselines
We compare the proposed DeTracker with a range of state-of-the-art trackers, including both TBD-based approaches, such as DeepSORT [44], CKDNet-SMTNet [12], and ByteTrack [59], and JDT-based approaches, including CenterTrack [66], FairMOT [60], DSFNet [49], MP2Net [63], DAF [54], PIFTrack [5], and MOSAIC-Tracker [69].
To ensure fair comparisons, all methods are trained and evaluated under the same experimental settings, with hyperparameters and training configurations adopted from the original papers or official open-source implementations.
@@ V-A3 Evaluation Metrics
For satellite video multi-object tracking, we adopt MOTA to measure overall tracking performance, particularly reflecting detection accuracy, and IDF1 to evaluate identity consistency, which reflects the quality of track association:
$\mathrm{MOTA}$ $=1-\frac{FN+FP+ID_{S}}{GT},$ (18)
@@
where $FP$, FNFN, and $ID_{S}$ denote false positives, false negatives, and identity switches, respectively; $GT$ is the number of ground-truth objects; $ID_{TP}$, $ID_{FP}$, and $ID_{FN}$ are the numbers of true-positive, false-positive, and false-negative identity matches. In addition, we report identity precision and recall (IDP and IDR), as well as standard MOT statistics including Precision (Prcn), Recall (Rcll), Mostly Tracked (MT), and Mostly Lost (ML), which provide complementary insight into tracking quality and robustness.
@@ V-A4 Implementation Details
Prior to training, standard data augmentation techniques were applied, including inversion, random flipping, and rotation. The initial learning rate was set to $1.25\times 10^{-4}$ and decayed using a cosine annealing schedule. The model was trained for 150 epochs with a batch size of 3. The loss weighting parameters $\alpha$ and $\beta$ were set to 0.2 and 0.3, respectively. The number of input frames $k$ was set to 5. All experiments were conducted on a single NVIDIA A100 Tensor Core GPU.
V-B Comparison with the State-of-the-Art
@@ V-B1 Quantitative Analysis
Table I presents the quantitative comparison on the SDM-Car-SU dataset. DeTracker consistently achieves the best overall performance across all motion intensity subsets (U1–U3), where U1–U3 correspond to increasing platform motion intensities from 1 to 3 pixels per frame. Specifically, in terms of the two core metrics, DeTracker achieves 61.1%, 55.3%, 52.4% in MOTA and 68.7%, 67.2%, 63.8% in IDF1 on U1, U2, and U3, respectively, outperforming all competing methods by a clear margin. Compared with the second-best method, PIFTrack [5], DeTracker yields an average improvement of 5.1% in MOTA and 2.9% in IDF1 across the three motion intensity levels. Additionally, it substantially reduces IDs, FP, and FN, indicating stronger robustness in data association and improved target recall in unstabilized videos.
To further evaluate generalization under real on-orbit conditions, we conduct experiments on real unstabilized satellite videos. As reported in Table II, DeTracker achieves 45.3% MOTA and 67.8% IDF1, surpassing PIFTrack by 8.3% and 7.7%, respectively. These results demonstrate that DeTracker not only performs well on simulated data but also maintains strong accuracy and robustness in real unstabilized satellite scenarios.
@@
FairMOT [60] DSFNet [49] MP2Net [63] PIFTrack [5] DeTracker (Ours) Ground truth
Figure 6: Examples of detection results for vehicles moving in different directions. From the first line to the third line are frames 50, 120, and 189 respectively, and vehicles undergo lateral and longitudinal motion. Green boxes denote true positives and correct predictions, red boxes indicate false positives, and yellow boxes represent false negatives.
@@
Input frame FairMOT [60] DSFNet [49] MP2Net [63] PIFTrack [5] DeTracker (Ours) Ground truth
@@ V-B2 Qualitative Analysis
We visualize detection results on randomized camera-shake video sequences with diverse vehicle motion directions for different methods, as illustrated in Fig. 6. Under complex backgrounds and severe platform jitter, most baseline methods suffer from frequent false positives and missed detections, particularly in regions with strong illumination changes, shadows, and road boundaries, where bounding boxes often drift or misalign. In contrast, DeTracker accurately localizes vehicle targets across consecutive frames, reducing both false positives and missed detections.
Fig. 7 further visualizes trajectory predictions from different trackers, where all trajectories are projected onto the last frame of each sequence and color coded by vehicle identity. As shown in the first row, DeTracker generates smooth and continuous trajectories that closely follow the ground truth paths, remain well aligned with road geometry, and exhibit minimal jitter. In contrast, most baselines suffer from notable trajectory fluctuations at turning regions, with some trajectories drifting along background structures. In the second row, baseline methods produce fragmented or discontinuous trajectories, indicating identity switches and unstable temporal association. Notably, MP2Net and DSFNet exhibit clear trajectory offsets and breaks, whereas DeTracker consistently predicts complete and coherent trajectories. We attribute this performance gain to the GLMD module, which decouples global background motion and enables feature-level alignment, and the TDFP module, which strengthens cross-frame dependency modeling, thereby preserving trajectory coherence and identity consistency over long-term tracking.